# A subset $S$ of vertices in a planar graph $G$ is a free set if, for every set $P$ of $|S|$ points in the plane, there exists a straight-line crossing-free drawing of $G$ in which vertices of $S$ are mapped to distinct points in $P$. In this survey, we review

### v1

@@ 3.1 Subclasses of planar graphs with linear-sized free sets
Note that the preceding argument holds even if the definition of level planar drawings is relaxed to allow: consecutive vertices within a level to be adjacent; edges between consecutive levels to be strictly $y$-monotone instead of straight; and even edges to go between non-consecutive levels (as long as they are not too far). This leads to the following definition: An $s$-span weakly level planar drawing is a crossing-free drawing in the plane, such that the vertices are placed on a sequence of parallel lines, where each edge $e$ is either a straight-line segment between two consecutive vertices on the same level (called, a horizontal edge, or a strictly $y$-monotone curve that intersects at most $s+1$ levels (called, a vertical edge). A graph is $s$-span weakly level planar if it has an $s$-span weakly level planar drawing. See [36] for example, for a similar definition. The following lemma formalizes the usefulness of this notion for obtaining free sets. This connection between $1$-span weakly level planar level planar drawings have been observed by [48].
@@
Refer to Fig. 4. Number the levels in an $s$-span weakly level planar drawing $\Gamma$ of $G$ by $0,1,2\dots$. Then for some $i\in\{0,1,\dots,s\}$, the union of levels $j=i\mod(s+1)$ has at least $\lceil n/(s+1)\rceil$ vertices, $S$ of $G$. Moreover $G[S]$ is a forest of paths (induced by the horizontal edges and the vertices on these levels). As in the case of level planar graphs, depict now a closed curve $C$ such that $C$ contains these levels of $\Gamma$ (and thus the vertices of $S$) such that $C$ intersects each vertical edge of $\Gamma$ at most once. This is possible since vertical edges are $y$-monotone and their endpoints lie on the levels whose difference (the absolute value) is at most $s$. Such a curve $C$ is a proper-good curve and thus by 1, $S$ a proper-good set whose vertices are ordered by their appearance on $C$. ∎
@@
Exactly the same proof used to show Lemma 2 shows the following “hereditary” variant:
@@
These lemmas have several immediate consequences. Firstly, as argued above, breath-first-search leveling of trees can be easily turned into level planar drawings (and thus, $1$-span weakly level planar drawings). Similarly, breath-first-search leveling of outerplanar graphs can be turned into $1$-span weakly level planar drawings of such graphs, as proved by Felsner et al. [32]. A natural way to draw the $n\times n$ grid graph is a 1-level planar drawing. More generally, Bannister et al. [7] show that squaregraphs are $1$-span weakly level planar. A squaregraph is a graph that has a crossing-free drawing in which each bounded face is a $4$-cycle and each vertex either belongs to the unbounded face or has four or more incident edges. The same authors also show that Halin graphs are $1$-span weakly level planar. Giacomo et al. [36] identified several classes of planar graphs that are $s$-span weakly level planar, for some constant $s$.
@@
Let $G$ be a $n$-vertex tree, outerplanar graph, Halin graph, or square graph and let $X$ be any subset of vertices of $G$. The $G$ has a free set $S\subseteq X$ of size at least $\lceil|X|/2\rceil$. In particular, $G$ has a free set of size at least $\lceil n/2\rceil$.
The $\lceil n/2\rceil$ lower bound for trees appears in [17] in the context of untangling. The extension to outerplanar graphs appears in [38], also in the context of untangling. The proof for outerplanar graphs by Goaoc et al. [38] produces a proper-good curve that is qualitatively different than the one in 1. See Fig. 5. In particular, the curve obtained there is contained in the closure of the outer face of $G$ and contains $G$ in the closure of its interior. This distinction turns out to be important when studying general planar graphs. Goaoc et al. [38] prove the following result although it is not stated in this form. We include a variant of their proof.
@@
Figure 5: An edge-maximal outerplane graph $G$; (b) a proper good curve guaranteed by Lemma 4; and (c) a proper good curve guaranteed by Lemma 2.
@@
Let $G$ be an edge-maximal $n$-vertex outerplane graph for some $n\geq 4$. Then there exists a proper-good curve $C$ that is contained in the closure of the outer face and $C$, that contains $G$ in the closure of its interior, and that contains at least $\frac{n}{2}+1$ vertices of $G$. Thus the vertices in $C$ form a proper-good set of size at least $\frac{n}{2}+1$.
Proof.
Since $G$ is edge-maximal and $n\geq 4$, the boundary of the outer face is a cycle, denoted by $O$. Consider a set $S\subseteq V(G)$ that has the following two properties: $G[S]$ induces a forest of paths and all the edges of $G[S]$ lie on $O$. It is easy to see that every such set $S$ is a proper-good set of $G$. Moreover, it is not hard to produce a proper-good curve that is contained in the closure of the outer face of $G$ and contains the vertices and edges of $G[S]$ by closely tracing the outer face of $G$.
We now prove that $G$ has such a set $S$ of claimed size. Let $T$ denote the weak dual of $G$. $T$ is a tree on at least two vertices since $n\geq 4$. Each vertex of $T$ has a degree $1$, $2$ or $3$, so $T$ is a binary tree. Let $t_{1}$ and $t_{3}$ denote the number of degree $1$ and degree $3$ vertices in $T$, respectively. It is well known that in every binary tree $t_{1}=t_{3}+2$.
Consider now the graph $G^{\prime}$ obtained from $G$ by removing the edges with both endpoints on the outer face (in other words, $G$’ contains only the chords of $O$). Every independent set in $G^{\prime}$ meets the two conditions imposed on $S$ earlier. Thus it remains to prove that $G^{\prime}$ has independent set $S$ of size $\frac{n}{2}+1$. Construct $S$ greedily in $G^{\prime}$ as follows: put in $S$ the vertex of $G^{\prime}$ of minimum degree; remove that vertex and its neighbors from $G^{\prime}$ to obtain a new $G^{\prime}$; and, repeat. $S$ is clearly an independent set in $G^{\prime}$. It remains to show that it has the claimed size. In the moment a vertex is placed in $S$ its degree in current $G^{\prime}$ was 0, 1 or 2. Let $n_{i}$ denote the number of vertices that had degree $i$ when they were placed in $S$. Thus $|S|=n_{0}+n_{1}+n_{2}$. From the description of the algorithm, it follows that $n=n_{0}+2n_{1}+3n_{2}$. Finally, each leaf of $T$ contributes $1$ to $n_{0}$. Thus $n_{0}\geq t_{1}$. Each vertex of $G^{\prime}$ that contributes to $n_{2}$ corresponds to a unique face whose dual vertex has degree $3$ in $T$. Thus $n_{2}\leq t_{3}$. Combining that inequality with the earlier obtained equality, $t_{1}=t_{3}+2$, we get that $n_{0}\geq n_{2}+2$. To summarize, we have the two equations and one inequality:
@@
Replacing $n_{1}$ in the first equality with $n_{1}=(n-n_{0}-3n_{2})/2$ obtained from the second equality gives $|S|=\frac{n}{2}+\frac{n_{0}-n_{2}}{2}$. Combined with the last inequality, we obtain the claimed result. ∎
Ravsky and Verbitsky [48] extend this result even further to show that all graphs of treewidth $2$ have free collinear sets of linear size. This extension does not follow from the argument used to prove Lemma 2, since there is a class of planar graphs of treewidth at most $2$ whose graphs do not have $s$-span weakly level planar drawings for any fixed $s$ [11].
Having a weakly level planar drawing of small span is a too strong of a condition to provide a general tool for finding large free sets in wider classes of planar graphs. In addition to the example of $2$-trees, there are planar graphs of bounded pathwidth, illustrated in Fig. 6 that do not have $o(n)$-span weakly level planar drawings (a result that can be derived from Biedl [11, Theorem 5]) and yet each graph in the class has an induced path of length $(n-1)/2$ that can be covered by a proper-good curve and thus the vertices on that path form a free set of size at least $(n-1)/2$, by 1.
@@
3 implies that all $n$-vertex planar graphs of treewidth $\Omega(\sqrt{n})$ have free set of linear size (a vast generalization of the above observation that square grids have linear size free sets).
Rather than consider planar graphs of small treewidth, one can also consider planar graphs of small (maximum) degree. Da Lozzo et al. [24] prove the following result in this vein:
@@
Every $n$-vertex planar triconnected cubic graph has a free set of size at least $\lceil\frac{n}{4}\rceil$
@@ 3.2 Free Sets in Planar Graphs
A version of 5 with an $\Omega(\sqrt{n})$ bound is attributed to Bose et al. [17] though the authors at the time were working on the untangling problem discussed in the introduction, so their result is never stated in terms of free sets. They prove that $\operatorname{fix}_{\mathcal{G}}(n)=\Omega(n^{1/4})$ by proving that a triangulation $G$ either contains an induced outerplane graph of size $\Omega(\sqrt{n})$ or a free-collinear set of size $\Omega(\sqrt{n})$. At the time, the equivalence between proper-good sets, collinear sets, and free-collinear sets was not known, so their proof includes both combinatorial and geometric elements (including a proof that free-collinear and free sets are equivalent). In the following, we extract these combinatorial elements from [17] and use Lemma 4 (proven in [38]) to give a self-contained proof of 5 with the best currently known bound of $\sqrt{n/2}$.
@@
Let $G$ be a near-triangulation, and let $v$ and $w$ be two points (possibly vertices) on the boundary of the outer face of $G$ such that $vw$ is not an edge of the outer face of $G$. Then there exists a simple open curve $C:(0,1)\to\operatorname{int}(G)$ with endpoints $v$ and $w$ that is proper with respect to $G$.
@@
(Da Lozzo et al. [24] give a 2-line proof of Lemma 5 using Tutte’s convex embedding theorem.) If $vw$ if a (internal) edge of $E(G)$ then the curve $C$ consists of the edge $vw$. Then $C$ contains the edge $vw$, intersects each other edge incident to $v$ or $w$ in one point, and does not intersect any other edge of $G$, so $C$ is proper with respect to $G$.
Refer to caption Refer to caption Refer to caption
Figure 7: The near-triangulation $G$, the curve $D$ from $f_{v}$ to $f_{w}$, and the proper-good curve $C$ from $v$ to $w$.
Since $vw$ is not an internal edge of $G$ and since $G$ is a near-triangulation, there is no internal face of $G$ that contains both $v$ and $w$ on its boundary. Refer to Fig. 7. Let $f_{v}$ be some internal face that contains $v$ and $f_{w}$ some internal face that contains $w$. Let $G^{+}$ be the weak dual of $G$. Let $D:[0,1]\to\operatorname{int}(G)$ be a simple open curve with one endpoint in $f_{v}$ and one in $f_{w}$ such that $D$ does not contain any vertices of $G$, as illustrated in Fig. 7(b). The sequence of faces of $G$ intersected by $D$ defines a walk $W$ in $G^{+}$ with endpoints $f_{v}$ and $f_{w}$. The walk $W$ contains a path $P$ in $G^{+}$ from $f_{v}$ to $f_{w}$. When traversing the path $P$ from $f_{v}$ to $f_{w}$, let $f_{v}^{\prime}$ be the last face on $P$ that contains $v$ on its boundary; and, let $f_{w}^{\prime}$ be the first face on $P$ that contains $w$ on its boundary. Let the resulting sub-path between $f_{v}^{\prime}$ and $f_{w}^{\prime}$ be denoted by $P^{\prime}$.
Let $e^{+}_{1},\ldots,e^{+}_{t}$ be the sequence of edges in $P^{\prime}$. For each $i\in\{1,\ldots,t\}$, the edge $e_{i}^{+}$ in $G^{+}$ corresponds to some inner edge $e_{i}$ of $G$. For each $i\in\{1,\ldots,t\}$, let $p_{i}$ be any point in the interior of $e_{i}$, let $p_{0}:=v$, and let $p_{t+1}=w$. Then, for each $i\in\{1,\ldots,t+1\}$ there is an inner face $f_{i}$ of $G$ with the points $p_{i-1}$ and $p_{i}$ on the boundary of $f_{i}$. Construct an open simple curve $C$ that visits $p_{0},\ldots,p_{t+1}$ in order in such a way that the portion of $C$ between $p_{i-1}$ and $p_{i}$ is contained in the interior of $f_{i}$, for each $i\in\{1,\ldots,t+1\}$. By the choice of $f_{v}^{\prime}$, no edge incident to $v$ intersects $C$ other than in $v$. Similarly, by the choice of $f_{w}^{\prime}$, no edge incident to $w$ intersects $C$ other than in $v$. Thus, the curve $C$ satisfies the requirements of the lemma. ∎
@@
If $G_{0}$ is a triangulation such that $G$ is a spanning subgraph of $G_{0}$, then clearly, any free set in $G_{0}$ is also a free set in $G$. Thus we may assume, without loss of generality, that $G$ is a triangulation.
The steps in this proof are illustrated in Fig. 8. Fix a crossing-free drawing of $G$ and let $v_{1}$, $v_{2}$, and $v_{n}$ be the three vertices on the outer face of $G$.
@@
We use a canonical ordering $v_{1},\ldots,v_{n}$ of $V(G)$, which has the following property: For each $i\in\{3,\ldots,n\}$, the induced graph $G_{i}:=G[\{v_{1},\ldots,v_{i}\}]$ is a near-triangulation that contains the vertex $v_{i}$ and the edge $v_{1}v_{2}$ on its outer face. The existence of this ordering is proven by, for example, de Fraysseix et al. [34]. In this ordering, the introduction of $v_{i}$ introduces $d_{i}\geq 2$ new edges $v_{i}w_{i,1},\ldots,v_{i}w_{i,d_{i}}$ to $G_{i}$ that do not appear in $G_{i-1}$. The edges $e_{i}:=v_{i}w_{i,1}$ and $e^{\prime}_{i}:=v_{i}w_{i,d_{i}}$ appear on the outer face of $G_{i}$ and the vertices $w_{i,2},\ldots,w_{i,d_{i}-1}$ on the outer face of $G_{i-1}$ no longer appear on the outer face of $G_{i}$, as illustrated in Fig. 8(a). This defines a sequence of frames $F_{1},\ldots,F_{n}$, where $V(F_{i}):=\{v_{1},\ldots,v_{i}\}$ and $E(F_{i}):=\bigcup_{j=1}^{i}\{e_{i},e^{\prime}_{i}\}$. We treat each frame $F_{i}$ as a directed acyclic graph with a single source $v_{1}$ and a single sink $v_{2}$ where the direction of edges $e_{i}$ and $e_{i}^{\prime}$ are the same in each frame $F_{i},\ldots,F_{n}$ that includes these two edges. [^7] For readers familiar with Schnyder Woods [52], the final frame $F_{n}$ can be obtained by taking the union of two trees $T_{1}$ and $T_{2}$ (rooted at $v_{1}$ and $v_{2}$, respectively) in a Schnyder Woods, directing each edge of $T_{1}$ away from its root and directing each edge of $T_{2}$ towards its root, as illustrated in Fig. 8(b). Since the final frame $F:=F_{n}$ is a directed acyclic graph, its transitive closure defines a partially ordered set $(V(G),\prec)$ in which $v\prec w$ if and only if $F$ contains a directed path from $v$ to $w$.
First, consider some maximal chain $x_{1}\prec\cdots\prec x_{k}$ in this partial order. Since $F$ is maximal, $x_{i}x_{i+1}$ is an edge of $F$ for each $i\in\{1,\ldots,k-1\}$. Since $F$ has a single source $v_{1}$ and a single sink $v_{2}$, $x_{1}=v_{1}$ and $x_{k}=v_{2}$. Thus, $C_{x}:=x_{1},\ldots,x_{k}$ is a cycle in $G$, as illustrated in Fig. 8(c).
@@
Suppose that $G$ contains some edge $x_{i}x_{j}$ that is a chord of $C_{x}$. We now argue that the edge $x_{i}x_{j}$ is embedded in the interior of the cycle $C_{x}$. Without loss of generality, suppose $x_{i}$ appears after $x_{j}$ in the canonical order, so $x_{i}=v_{a}$ and $x_{j}=v_{b}$ for some $a>b$. Then the edge $x_{i}x_{j}=v_{a}v_{b}$ is in the graph $G_{a}=G[\{v_{1},\ldots,v_{a}\}]$, so $v_{b}\in\{w_{a,1},\ldots,w_{a,d_{a}}\}$. The two neighbours $x_{i-1}$ and $x_{i+1}$ of $x_{i}$ in $C_{x}$ are not in the set $w_{a,2},\ldots,w_{a,d_{a}-1}$, so the interior of the cycle $C_{x}$ contains the interior of the cycle $v_{a},w_{a,1},\ldots,w_{a,d_{a}}$. Therefore the edge $v_{a}v_{b}=x_{i}x_{j}$ is in the interior of $C_{x}$. (Indeed, the only possibility is that $v_{b}\in\{w_{a,1},w_{a,d_{a}}\}$ )

### v2

@@ 3.1 Subclasses of planar graphs with linear-size free sets
Note that the preceding argument holds even if the definition of level planar drawings is relaxed to allow: consecutive vertices within a level to be adjacent; edges to go between non-consecutive levels, as long as they are not too far apart; and edges between different levels to be strictly $y$-monotone instead of straight. This leads to the following definition by Bekos et al. [10]: An $s$-span weakly level planar drawing is a crossing-free drawing in the plane, such that the vertices are placed on a sequence of parallel lines, where each edge $e$ is either a straight-line segment between two consecutive vertices on the same level (called a horizontal edge) or a strictly $y$-monotone curve that intersects at most $s+1$ levels (called a vertical edge). A graph is $s$-span weakly level planar if it has an $s$-span weakly level planar drawing. The following lemma formalizes the usefulness of this notion for obtaining free sets. This connection between $1$-span weakly level planar level planar drawings have been observed by [50].
@@
Refer to Fig. 4. Number the levels in an $s$-span weakly level planar drawing $\Gamma$ of $G$ by $0,1,2\dots$. Then for some $i\in\{0,1,\dots,s\}$, the union of levels $j=i\mod(s+1)$ has at least $\lceil n/(s+1)\rceil$ vertices, $S$ of $G$. Moreover, $G[S]$ is a forest of paths (induced by the horizontal edges and the vertices on these levels). As in the case of level planar graphs, construct a closed curve $C$ that contains these levels of $\Gamma$ (and thus the vertices of $S$) and such that $C$ intersects each vertical edge of $\Gamma$ at most once. This is possible since vertical edges are $y$-monotone and their endpoints lie on the levels whose difference (the absolute value) is at most $s$. Such a curve $C$ is a proper-good curve and thus by 1, $S$ is a free set whose vertices are ordered by their appearance on $C$. ∎
@@
Exactly the same proof shows the following “hereditary” variant of Lemma 2:
@@
These lemmas have several immediate consequences. Firstly, as argued above, breath-first-search leveling of trees can be easily turned into level planar drawings (and thus $1$-span weakly level planar drawings). Similarly, breath-first-search leveling of outerplanar graphs can be turned into $1$-span weakly level planar drawings of such graphs, as proved by Felsner et al. [32]. The natural way to draw the $n\times n$ grid graph is a 1-level planar drawing. More generally, Bannister et al. [6] show that squaregraphs are $1$-span weakly level planar. A squaregraph is a graph that has a crossing-free drawing in which each bounded face is a $4$-cycle and each vertex either belongs to the unbounded face or has four or more incident edges. The same authors also show that Halin graphs are $1$-span weakly level planar. Di Giacomo et al. [36] identified several classes of planar graphs that are $s$-span weakly level planar, for some constant $s$.
@@
Let $G$ be a $n$-vertex tree, outerplanar graph, Halin graph, or square graph and let $X$ be any subset of the vertices of $G$. The $G$ has a free set $S\subseteq X$ of size at least $\lceil|X|/2\rceil$. In particular, $G$ has a free set of size at least $\lceil n/2\rceil$.
The $\lceil n/2\rceil$ lower bound for trees appears in [17] in the context of untangling. The extension to outerplanar graphs appears in [39], also in the context of untangling. For outerplanar graphs, a different construction of a proper-good curve was given by Goaoc et al. [39]. It has a special structure, and the bound is slightly stronger. See Fig. 5(c). In particular, the curve obtained there is contained in the closure of the outer face of $G$ and contains $G$ in the closure of its interior. This structure turns out to be important when studying general planar graphs. Goaoc et al. [39] prove the following result, although it is not stated in this form. We include a variant of their proof.
@@
Figure 5: An edge-maximal outerplane graph $G$; (b) a proper-good curve guaranteed by Lemma 2/1; and (c) a proper-good curve guaranteed by Lemma 4.
@@
Let $G$ be an $n$-vertex biconnected outerplane graph for some $n\geq 4$. Then there exists a proper-good curve $C$ that is contained in the closure of the outer face of $G$, that contains $G$ in the closure of its interior, and that passes through at least $\lceil n/2\rceil+1$ vertices of $G$. Thus the vertices in $C$ form a proper-good set of size at least $\lceil n/2\rceil+1$.
Before proving this lemma, we emphasize that the condition that $C$ is contained in the closure of the outer face of $G$ does not imply the second condition—namely, that $G$ lies in the interior of $C$. These two properties will be important for identifying large free sets in general planar graphs. In particular, they ensure that $C$ can closely follow the outer face of an induced outerplane subgraph of a planar graph.
Proof of Lemma 4.
Since $G$ is biconnected and $n\geq 4$, the boundary of the outer face is a cycle, denoted by $O$. We may assume that $G$ is edge-maximal so that the boundary of each inner face of $G$ is a cycle of length $3$. Consider a set $S\subseteq V(G)$ that has the following two properties: $G[S]$ induces a forest of paths and all the edges of $G[S]$ lie on $O$. It is easy to see that every such set $S$ is a proper-good set of $G$. Moreover, it is not hard to produce a proper-good curve that is contained in the closure of the outer face of $G$ and contains the vertices and edges of $G[S]$ by closely tracing the outer face of $G$.
We now prove that $G$ has such a set $S$ of claimed size. Let $T$ denote the weak dual of $G$. Then $T$ is a tree on at least two vertices since $n\geq 4$. Each vertex of $T$ has degree $1$, $2$ or $3$, so $T$ is a binary tree. Let $t_{1}$ and $t_{3}$ denote the number of degree $1$ and degree $3$ vertices in $T$, respectively. It is well known that in every binary tree $t_{1}=t_{3}+2$.
Consider now the graph $G^{\prime}$ obtained from $G$ by removing the edges on the outer face (in other words, $G^{\prime}$ contains only the chords of $O$). Every independent set in $G^{\prime}$ meets the two conditions imposed on $S$ earlier. Thus it remains to prove that $G^{\prime}$ has an independent set $S$ of size at least $\frac{n}{2}+1$. Construct $S$ greedily in $G^{\prime}$ as follows: put in $S$ the vertex of $G^{\prime}$ of minimum degree; remove that vertex and its neighbours from $G^{\prime}$ to obtain a new $G^{\prime}$; and repeat. $S$ is clearly an independent set in $G^{\prime}$. It remains to show that it has the claimed size. In the moment a vertex is placed in $S$, its degree in the current $G^{\prime}$ was 0, 1 or 2. Let $n_{i}$ denote the number of vertices that had degree $i$ when they were placed in $S$. Thus $|S|=n_{0}+n_{1}+n_{2}$. From the description of the algorithm, it follows that $n=n_{0}+2n_{1}+3n_{2}$. Finally, each leaf of $T$ contributes $1$ to $n_{0}$. Thus $n_{0}\geq t_{1}$. Each vertex of $G^{\prime}$ that contributes to $n_{2}$ corresponds to a unique face whose dual vertex has degree $3$ in $T$. Thus $n_{2}\leq t_{3}$. Combining that inequality with the earlier obtained equality, $t_{1}=t_{3}+2$, we get that $n_{0}\geq n_{2}+2$. To summarize, we have the two equations and one inequality:
@@
Replacing $n_{1}$ in the first equality with $n_{1}=(n-n_{0}-3n_{2})/2$ obtained from the second equality gives $|S|=\frac{n}{2}+\frac{n_{0}-n_{2}}{2}$. Combined with the last inequality, we obtain $|S|\geq n/2+1$, and the result follows from the fact that $|S|$ is an integer. ∎
Ravsky and Verbitsky [50] extend this result even further to show that all graphs of treewidth $2$ have free-collinear sets of linear size. This extension does not follow from the argument used to prove Lemma 2, since there is a class of planar graphs of treewidth at most $2$ whose graphs do not have $s$-span weakly level planar drawings for any fixed $s$ [11].
Having a weakly level planar drawing of small span is too strong a condition to provide a general tool for finding large free sets in wider classes of planar graphs. In addition to the example of $2$-trees, there are planar graphs of bounded pathwidth, illustrated in Fig. 6 that do not have $o(n)$-span weakly level planar drawings (a result that can be derived from Biedl [11, Theorem 5]) and yet each graph in the class has an induced path of length $(n-1)/2$ that can be covered by a proper-good curve and thus the vertices on that path form a free set of size at least $(n-1)/2$, by 1.
@@
3 implies that all $n$-vertex planar graphs of treewidth $\Omega(\sqrt{n})$ have a free set of linear size (a vast generalization of the above observation that square grids have linear size free sets).
Rather than considering planar graphs of small treewidth, one can also consider planar graphs of small (maximum) degree. Da Lozzo et al. [24] prove the following result in this vein:
@@
Every $n$-vertex planar triconnected cubic graph has a free set of size at least $\lceil\frac{n}{4}\rceil$.
@@ 3.2 Free Sets in Planar Graphs
A version of 5 with an $\Omega(\sqrt{n})$ bound is due to Bose et al. [17] though the authors at the time were working on the untangling problem discussed in the introduction, so their result is never stated in terms of free sets. They prove that $\operatorname{fix}_{\mathcal{G}}(n)=\Omega(n^{1/4})$ by proving that a triangulation $G$ either contains an induced outerplane graph of size $\Omega(\sqrt{n})$ or a free-collinear set of size $\Omega(\sqrt{n})$. At the time, the equivalence between proper-good sets, collinear sets, and free-collinear sets was not known, so their proof includes both combinatorial and geometric elements (including a proof that free-collinear and free sets are equivalent). In the following, we extract these combinatorial elements from [17] and use Lemma 4 (proven in [39]) to give a self-contained proof of 5 with the best currently known bound of $\sqrt{n/2}$.
@@
Let $G$ be a near-triangulation, and let $v$ and $w$ be two points (possibly vertices) on the boundary of the outer face of $G$ such that no edge of the outer face of $G$ contains $v$ and $w$. Then there exists a simple curve $C:[0,1]\to\operatorname{int}(G)\cup\{v,w\}$ with endpoints $v$ and $w$ that is proper with respect to $G$.
@@
If $vw$ is an (internal) edge of $E(G)$, then the curve $C$ consists of the edge $vw$, which clearly satisfies the requirements of the lemma.
Figure 7: Constructing a proper curve $C$ from $v$ to $w$ through $\operatorname{int}(G)$.
Otherwise, since the edge $vw$ is not an edge of $G$ and since $G$ is a near-triangulation, there is no internal face of $G$ that contains both $v$ and $w$ on its boundary. Refer to Fig. 7. Let $G^{+}$ be the weak dual of $G$. Let $P$ be a shortest path in $G^{+}$ having one endpoint in a face that contains $v$ and the other endpoint in a face that contains $w$. Let $e^{+}_{1},\ldots,e^{+}_{t}$ be the sequence of edges in $P$. For each $i\in\{1,\ldots,t\}$, the edge $e_{i}^{+}$ in $G^{+}$ corresponds to some inner edge $e_{i}$ of $G$. For each $i\in\{1,\ldots,t\}$, let $p_{i}$ be any point in the interior of $e_{i}$, let $p_{0}:=v$, and let $p_{t+1}=w$. Then, for each $i\in\{1,\ldots,t+1\}$ there is an inner face $f_{i}$ of $G$ with the points $p_{i-1}$ and $p_{i}$ on the boundary of $f_{i}$. Construct an open simple curve $C$ that visits $p_{0},\ldots,p_{t+1}$ in order in such a way that the portion of $C$ between $p_{i-1}$ and $p_{i}$ is contained in the interior of $f_{i}$, for each $i\in\{1,\ldots,t+1\}$.
Clearly $C$ is contained in $\operatorname{int}(G)\cup\{v,w\}$, so all that remains is to show that $C$ is proper with respect to $G$. By construction, $C$ intersects each of the edges $e_{1},\ldots,e_{t}$, $C$ intersects the edges of $G$ incident to $v$, $C$ intersects the edges of $G$ incident to $w$, and $C$ avoids all other edges of $G$. Since $P$ is a shortest path, none of $e_{1},\ldots,e_{t}$ is incident to $v$ or $w$. Since $P$ is a path, $C$ intersects each of $e_{1},\ldots,e_{t}$ in a single point. Finally, $C$ intersects each of the edges of $G$ incident to $v$ (respectively $w$) in a single point, namely $v$ (respectively, $w$). Thus $C$ is proper with respect to $G$. ∎
@@
If $G_{0}$ is a triangulation such that $G$ is a spanning subgraph of $G_{0}$, then clearly, any free set in $G_{0}$ is also a free set in $G$. Thus we may assume, without loss of generality, that $G$ is a triangulation. The steps in this proof are illustrated in Fig. 8. Fix a crossing-free drawing of $G$ and let $v_{1}$, $v_{2}$, and $v_{n}$ be the three vertices on the outer face of $G$.
@@
We use a canonical ordering $v_{1},\ldots,v_{n}$ of $V(G)$, which has the following property: For each $i\in\{3,\ldots,n\}$, the induced graph $G_{i}:=G[\{v_{1},\ldots,v_{i}\}]$ is a near-triangulation that contains the vertex $v_{i}$ and the edge $v_{1}v_{2}$ on its outer face. The existence of such an ordering is proven by de Fraysseix et al. [34]. In this ordering, the introduction of $v_{i}$ introduces $d_{i}\geq 2$ new edges $v_{i}w_{i,1},\ldots,v_{i}w_{i,d_{i}}$ to $G_{i}$ that do not appear in $G_{i-1}$. The edges $e_{i}:=v_{i}w_{i,1}$ and $e^{\prime}_{i}:=v_{i}w_{i,d_{i}}$ appear on the outer face of $G_{i}$ and the vertices $w_{i,2},\ldots,w_{i,d_{i}-1}$ on the outer face of $G_{i-1}$ no longer appear on the outer face of $G_{i}$, as illustrated in Fig. 8(a). This defines a frame $F$ where $V(F):=V(G)$ and $E(F):=\{v_{1}v_{2}\}\cup\bigcup_{j=3}^{n}\{e_{i},e^{\prime}_{i}\}$. We treat $F$ as a directed acyclic graph where $v_{1}v_{2}$ is directed towards $v_{2}$ and $w_{i,1},v_{i},w_{i,d_{i}}$ is a directed path. In this way, $F$ has a single source, $v_{1}$ and a single sink, $v_{2}$. [^12] For readers familiar with Schnyder woods [53], $F$ can be obtained by taking the union of two trees $T_{1}$ and $T_{2}$ (rooted at $v_{1}$ and $v_{2}$, respectively) in a Schnyder Wood, directing each edge of $T_{1}$ away from its root and directing each edge of $T_{2}$ towards its root, as illustrated in Fig. 8(b). Since the final frame $F$ is a directed acyclic graph, its transitive closure defines a partially ordered set $(V(G),\prec)$ in which $v\prec w$ if and only if $F$ contains a directed path from $v$ to $w$.
@@
First, consider some maximal chain $x_{1}\prec\cdots\prec x_{k}$ in this partial order. Since $F$ is maximal, $x_{i}x_{i+1}$ is an edge of $F$ for each $i\in\{1,\ldots,k-1\}$. Since $F$ has a single source $v_{1}$ and a single sink $v_{2}$, $x_{1}=v_{1}$ and $x_{k}=v_{2}$. Thus $C_{x}:=x_{1},\ldots,x_{k}$ is a cycle in $G$, as illustrated in Fig. 8(c). Suppose that $G$ contains some edge $x_{i}x_{j}$ that is a chord of $C_{x}$. We now argue that the edge $x_{i}x_{j}$ is embedded in the interior of the cycle $C_{x}$. Without loss of generality, suppose $x_{i}$ appears after $x_{j}$ in the canonical order, so $x_{i}=v_{a}$ and $x_{j}=v_{b}$ for some $a>b$. Then the edge $x_{i}x_{j}=v_{a}v_{b}$ is in the graph $G_{a}=G[\{v_{1},\ldots,v_{a}\}]$, so $v_{b}\in\{w_{a,1},\ldots,w_{a,d_{a}}\}$. The two neighbours $x_{i-1}$ and $x_{i+1}$ of $x_{i}$ in $C_{x}$ are not in the set $w_{a,2},\ldots,w_{a,d_{a}-1}$, so the interior of the cycle $C_{x}$ contains the interior of the cycle $v_{a},w_{a,1},\ldots,w_{a,d_{a}}$. Therefore, the edge $v_{a}v_{b}=x_{i}x_{j}$ is in the interior of $C_{x}$. (Indeed, the only possibility is that $v_{b}\in\{w_{a,1},w_{a,d_{a}}\}$ )